OpenStudy (anonymous):

PLZ HELP WILL MEDAL AND FAN What is the exact diameter of an area rug with a circumference of thirty-six feet? eighteen pi feet eighteen divided by pi feet thirty-six divided by pi feet thirty-six pi feet

OpenStudy (marihelenh):

\[C=(\pi)d\]

OpenStudy (marihelenh):

It is also equal to \[C=(\pi)r ^{2}\] So, if you put 36 in for the circumference, you get \[36=(\pi)r ^{2}\]

OpenStudy (anonymous):

thank you

OpenStudy (marihelenh):

From there you can square root both sides and get your answer.

OpenStudy (marihelenh):

Oh my goodness wait.

OpenStudy (anonymous):

mari you're wrong C=pi*d is the circumference while pi*r^2 is the area

OpenStudy (marihelenh):

I did that incorrectly. I realized what I did @RicardoDelfin So, @joannacaro123 the other equation is \[C=(\pi)2r\]

OpenStudy (anonymous):

so I multiply

OpenStudy (marihelenh):

You could put back 36 in for C. From there you would divide and get what r is equal to.

OpenStudy (anonymous):

so my answer is \[\frac{ 36 }{ } \]

OpenStudy (anonymous):



OpenStudy (marihelenh):